Find an explicit solution of the given initial-value problem

dy/dx = ye^(-x^2), y(5) = 1
1
Expert's answer
2020-07-27T18:55:46-0400

"\frac{dy}{dx} = ye^{-x^2}"

"\frac{dy}{y} = e^{-x^2}dx"

"\int\frac{du}{y}=\int e{^{-x^2}}dx"

But "\int e^{-x^2}dx"  is a non elementary integral. We can express it as a power series

"e^{-x^2}= 1 - \frac{x^2}{1!}+\frac{x^4}{2!}-\frac{x^6}{3!} +\frac{x^8}{4!}-.."

We can approximate it as

"e^{-x^2}= 1 - \frac{x^2}{1!}+\frac{x^4}{2!}-\frac{x^6}{3!}"

So differential equation transforms to

"\int\frac{du}{y}=\int" "(1 - \frac{x^2}{1!}+\frac{x^4}{2!}-\frac{x^6}{3!})dx"

=>ln|y| = "x - \frac{x^3}{3}+\frac{x^5}{10}-\frac{x^7}{42} + C"

By initial condition y(5)=1

0 = "5 - \frac{5^3}{3}+\frac{5^5}{10}-\frac{5^7}{42} + C"

C = 1584.286

So ln|y| = "x - \frac{x^3}{3}+\frac{x^5}{10}-\frac{x^7}{42} + 1584.286"

=> y = "e^{(x - \frac{x^3}{3}+\frac{x^5}{10}-\frac{x^7}{42} + 1584.286)}"

This is the explicit solution of the given differential equation


"\mathbf ALTERNATIVELY"

"\frac{dy}{dx} = ye^{-x^2}"

=> "\frac{dy}{y} = e^{-x^2}dx"

=> "\int\frac{du}{y}=\int e{^{-x^2}}dx"

As "\int e{^{-x^2}}dx" is not an elementary integral, we will express it as error function

=> ln|y| = ("\frac{\sqrtπ}{2}" ) erf(x) + C1, C1 is integration constant

Where erf(x) = "\int_0^x e{^{-t^2}}dt"

So y = "e^{ (\frac{\sqrtπ}{2}) erf(x) + C_1}"

=> y = C"e^{ (\frac{\sqrtπ}{2}) erf(x)}"

By initial condition y (5)=1

1 = C"e^{ (\frac{\sqrtπ}{2}) erf(5) }"

C = "e^{ -(\frac{\sqrtπ}{2}) erf(x) }"

So y = "e^{ (\frac{\sqrtπ}{2}) erf(x) - (\frac{\sqrtπ}{2}) erf(5) }"

=> y = "e^{ (\frac{\sqrtπ}{2}) [erf(x) - erf(5) ]}"

This is the explicit solution of the given differential equation